A baseball pitcher throws a fastball at a speed of 40.25 m/s (90. mph). The acceleration occurs as the pitcher holds the ball in his hand and moves through an almost straight-line distance of 3.40 m. Assuming that the acceleration is uniform, determine the acceleration. Compare the acceleration to the acceleration due to gravity (ratio or ‘G force’). 


1
Expert's answer
2021-02-22T10:23:03-0500

Answer

Accelaration is given by

"a=\frac{v_f^2-v_i^2}{2d}=\frac{(40.25)^2-0}{2\times3.4}\\=236.24m/s^2"


Now compare to gravity

"\frac{a}{g}=\frac{236.24}{9.8}=24.31"